OpenStudy (anonymous):

Antonio is seven less than four times his daughter's age. Antonio is 37; how old is his daughter? 11 12 18 6

OpenStudy (anonymous):

I just did one like this for you .. try to apply the same principles. Remember, "is" means equal sign.

OpenStudy (anonymous):

Antonio(37)=4*Age(x)-7 37=4A-7 +7 +7 44=4a /4 /4 A=11

OpenStudy (anonymous):

4x-7=37 4x=44 x=11